# Burgers equation from non-thermal stationary states in nearly-integrable gases

### v1

@@ V.3 Stochastic velocity swap models
where $\tilde{A}$ is some positive constant. From (81) we get the stationary state (39). The constant $A=\frac{\tilde{A}}{2\pi}$ is fixed by the condition $\varrho=\int{\rm d}\lambda\,\rho_{\mathrm{p}}^{\mathrm{st}}(\lambda)$. It follows that for a fixed state $\rho_{\mathrm{p}}^{(b)}$, $A$ is a function of $\varrho$ only and so $\frac{\partial\rho_{\mathrm{p}}^{\mathrm{st}}(\lambda,\varrho)}{\partial\varrho}=\frac{\partial\rho_{\mathrm{p}}^{\mathrm{st}}(\lambda,A)}{\partial A}\frac{{\rm d}A}{{\rm d}\varrho}$. By a direct computation we get
$\frac{\partial\rho_{\mathrm{p}}^{\mathrm{st}}(\lambda,A)}{\partial A}=-A^{-1}\rho_{\mathrm{p}}^{\mathrm{st}}(\lambda)(1-\chi n^{\mathrm{st}}(\lambda))$ (84)
@@
$\frac{\delta\mathcal{I}}{\delta\rho}\Big|_{\rho_{\mathrm{p}}^{\mathrm{st}}}\cdot a=\frac{(\lambda-u)\rho^{\mathrm{st}}(\lambda)(1-\chi n^{\mathrm{st}}(\lambda))}{\left(h_{0}|h_{0}\right)_{\mathrm{st}}}.$ (89)
@@
$\times\left[\delta\epsilon(\lambda)-\delta\epsilon(\lambda^{\prime})\right],$ (91)
where we used that for a stationary state equation (83) is satisfied. From the definition (81) it follows that $\delta\rho_{\mathrm{p}}=-\rho(1-\chi n)\delta\epsilon$, which can be checked by a direct calculation. Introducing
$b(\lambda)=-\gamma\frac{\left(h_{0}|h_{0}\right)_{\mathrm{st}}\,a(\lambda)}{\rho^{\mathrm{st}}(\lambda)(1-\chi n^{\mathrm{st}}(\lambda))}$ (92)
@@
Figure 4: Advective coefficient and diffusion coefficient as a function of $\varrho$ for the fermionic system coupled to the classical bath in the same state and $\omega$ as in Fig. 5
Figure 5: Molecular dynamics simulations for velocity swap models ($\gamma=0.08$, $\varrho^{(b)}=50.0$, $N=1600$) with initial states given by (104). We observe an excellent agreement between numerical simulations (histograms) and hydrodynamic equation (2) derived through Chapman-Enskog method (solid lines) for all cases. For symmetric states in bath system, both systems are described by diffusion equation (1). Upon introducing the asymmetry, the classical system is described by boosted diffusion equation. On the other hand, the fermionic model exhibits Burgers equation dynamics, as visible in the dynamically generated asymmetry in the initially Gaussian shape. Results from microscopic simulations are averaged over 1000 realizations.
@@
For the classical particles $\xi=0$ and the relation between the reference state and a stationary state is especially simple and given by (20). The convection velocity function $u(\varrho)$ simplifies to
@@ VI Numerical simulations
$\rho_{p}^{(b)}\propto e^{-(\lambda+\lambda_{B})^{2}/\sigma^{2}}+be^{-(\lambda-\lambda_{B})^{2}/\sigma^{2}}$ (104)
with $\sigma=6$, $\lambda_{B}=5.9$ and $b=1$ for symmetric states and $b=0.3$ for asymmetric states. The particles propagate freely and occasionally undergo two-body collisions with the bath. A collision amounts to velocity swap between system and a bath. When particles are classical and collisions happen at a constant rate the corresponding Master equation is precisely (37). Since the rate does not depend on the state of the system, such collisions can be efficiently implemented with standard Gillespie algorithm. The more complicated case is the simulation of fermionic kinetic equation given by (38). There, we partition the system into fluid cells and propose an algorithm, which effectively incorporates the Pauli principle into the collisions. Algorithms for both cases are presented in Appendix B.
The results are presented in Fig. 5. A very good agreement is found between molecular simulations and hydrodynamic equation (2). Depending on the symmetry of the state and particle statistics, we find both diffusion equation (1) and Burgers equation (2) for fermionic model with asymmetric state.
@@ VII Summary and discussion
In this work, we have considered a canonical problem of diffusion in which a minority species undergoes an irreversible motion due to coupling to the majority species whose dynamics is not monitored and which, as is customary to assume, is in a spatially homogenous state. Due to the conservation of particle number of the minority species, its density obeys a continuity equation. In the standard situation of classical particles this continuity equation leads to a (possibly non-linear) diffusion equation. In this work we identified the conditions under which the density follows instead the Burgers’ equation.
@@
The two sufficient conditions are: a) athermal and asymmetric state of the majority species (bath) and b) quantum statistics of minority species (system).
@@
We have demonstrated that these two conditions are fulfilled for effective dynamics of nearly integrable coupled gases. In such scenario, the integrability naturally makes athermal states possible. We have provided a kinetic explanation for appearance of such athermal states through limitations of two-body scattering in 1D. In our approach, we assume a large density imbalance between the two gases allowing for a separation of timescales between the slow dynamics in the dense tube (bath) and fast dynamics in the dilute tube (system). Starting from the GHD-Boltzmann description, we have derived an effective hydrodynamic equation for the density of the system particles (minority species). We have shown that the resulting equations become a standard diffusion equation when the bath is in a thermal state. Instead, for athermal state of a bath we obtain Burgers’ equation. Our method provides also expressions for transport coefficients: convective coefficient and diffusion constant.
@@
To unravel the essential ingredients leading to the Burgers’ equation we have considered also a simple stochastic model of two-particle scatterings between system and the bath. We observed an excellent agreement between numerical simulations and hydrodynamics derived from kinetic equation describing such random scattering events.
@@
Lastly, let us discuss the nature of the hydrodynamic density-density correlation function. Under the presence of ballistic mode with a nonlinear current the correlation function is expected to exhibit a breakdown of diffusive behavior at the largest scales [50]. Close to the integrability, we generically expect large diffusion constants, which push away the anomalous broadening to very large space-time scales. In this sense, we expect a scenario similar to the one analyzed in the context of Navier-Stokes equations [41]. We also note that one has to be careful when looking at asymptotically large scales. There are three-particle collision processes, which provide a mechanism for thermalization at late times, assuming that the density of the bath is finite. This leads ultimately to a breakdown of Burgers equation.
@@ Acknowledgements.
Appendix A Derivation of collision operator from GBBGKY hierarchy
For the purpose of presentation in this section we change the notation slightly. The quasimomentum dependence will be denoted in lower index and we introduce rotation matrices $R=1-\mathcal{T}n$. We also denote $\mathcal{A}_{\lambda_{1},\lambda_{2}}=R^{-t}_{\lambda_{1},\gamma}v_{\gamma}R^{t}_{\gamma,\lambda_{2}}$ and $\mathcal{C}_{\lambda_{1},\lambda_{2}}=R^{-t}_{\lambda_{1},\gamma}\rho_{\gamma}f_{\gamma}R^{-1}_{\gamma,\lambda_{2}}$. The upper indices $(i)$ denote tube index. We start with the following set of equations:
@@ Appendix A Derivation of collision operator from GBBGKY hierarchy
$g_{\lambda_{1},\lambda_{2}}=(R^{(1)})^{-t}_{\lambda_{1},u_{1}}(R^{(2)})^{-t}_{\lambda_{2},u_{2}}\bar{g}_{\gamma_{1},\gamma_{2}},$ (110)
@@
$G_{\lambda_{1},\lambda_{2}}(z,z^{\prime};t,t^{\prime})=\frac{1}{2\pi}\int{\rm d}k\,e^{ik(z+z^{\prime}-v_{\lambda_{1},\lambda_{2}}t)}\theta_{H}(t-t^{\prime})$ (112)
@@
$\bar{g}_{\lambda_{1},\lambda_{2}}(z,t)=\int{\rm d}z^{\prime}{\rm d}t^{\prime}G_{\lambda_{1},\lambda_{2}}(z,z^{\prime};t,t^{\prime})\bar{\mathcal{S}}_{\lambda_{1},\lambda_{2}}(z^{\prime},t^{\prime}).$ (113)
@@
In the next step we plug this expression into (107), replace the time argument of $\mathcal{S}_{\eta_{1},\eta_{2}}[\rho](z^{\prime},t-\tau)$ with $t$ and extend the integration range over $\tau$ to infinity. We also insert momentum representations of interaction potential $V(z)=\int{\rm d}ke^{ikz}\hat{V}(k)$ and use that $\int{\rm d}\lambda R^{-t}_{\lambda,\gamma}=2\pi\rho^{\rm t}_{\gamma}$. After these steps we find
$\partial_{t}\rho_{\lambda_{1}}^{(1)}=\mathcal{I}(\lambda_{1}),\qquad\mathcal{I}(\lambda_{1})=\partial_{\lambda_{1}}\left[(R^{(1)})^{-\rm{t}}_{\lambda_{1},\eta}\mathcal{I}_{0}(\eta)\right],$ (115)
@@
$\mathcal{I}_{0}(\lambda_{1})=2\pi\int{\rm d}kk^{2}\hat{V}^{2}(k)\int_{0}^{\infty}{\rm d}\tau\int{\rm d}\lambda_{2}e^{-ikv_{\lambda_{1}\lambda_{2}}\tau}\rho^{\rm t,(2)}_{\lambda_{2}}(R^{(1)})_{\lambda_{1},\eta_{1}}^{t}(R^{(2)})_{\lambda_{2},\eta_{2}}^{t}\int{\rm d}\eta_{3}\left(\partial_{\eta_{1}}\rho_{\eta_{1}}^{(1)}\mathcal{C}^{(2)}_{\eta_{2}\eta_{3}}-\partial_{\eta_{2}}\rho_{\eta_{2}}^{(2)}\mathcal{C}^{(1)}_{\eta_{1}\eta_{3}}\right)$. (116)
We proceed by using $\int_{0}^{\infty}{\rm d}\tau e^{i\tau x}=\pi\delta(x)+i\mathcal{P}(\frac{1}{x})$, where the second term is the principal value, which vanishes as the integrand is a symmetric function in $k$. In what follows we also use that $\int{\rm d}\lambda_{2}\,\mathcal{C}_{\lambda_{1},\lambda_{2}}=2\pi R_{\lambda_{1},\gamma}^{-t}\rho_{\gamma}f_{\gamma}\rho^{\rm t}_{\gamma}$ and $R^{t}_{\lambda,\gamma}\partial_{\gamma}\rho_{\gamma}=-\rho_{\lambda}f_{\lambda}\partial_{\lambda}\epsilon_{\lambda}$. We get
$\mathcal{I}_{0}(\lambda_{1})=-(2\pi)^{2}\pi\int{\rm d}k|k|\hat{V}^{2}(k)\int{\rm d}\lambda_{2}\delta(v_{\lambda_{1},\lambda_{2}})(\rho^{\rm t,(2)}_{\lambda_{2}})^{2}\rho_{\lambda_{1}}^{(1)}f_{\lambda_{1}}^{(1)}\rho_{\lambda_{2}}^{(2)}f_{\lambda_{2}}^{(2)}\left(\epsilon^{\prime}_{\lambda_{1}}-\frac{\rho^{\rm t}_{\lambda_{1}}}{\rho^{\rm t}_{\lambda_{2}}}\epsilon^{\prime}_{\lambda_{2}}\right).$ (117)
@@
Appendix B Molecular dynamics simulation algorithm

### v2

@@ V.3 Stochastic velocity swap models
where $\tilde{A}$ is some positive constant. From (82) we recover the stationary states (40) and (20). The constant $A=\frac{\tilde{A}}{2\pi}$ is fixed by the condition $\varrho=\int{\rm d}\lambda\,\rho_{\mathrm{p}}^{\mathrm{st}}(\lambda)$. It follows that for a fixed state $\rho_{\mathrm{p}}^{(b)}$, $A$ is a function of $\varrho$ only and so $\frac{\partial\rho_{\mathrm{p}}^{\mathrm{st}}(\lambda,\varrho)}{\partial\varrho}=\frac{\partial\rho_{\mathrm{p}}^{\mathrm{st}}(\lambda,A)}{\partial A}\frac{{\rm d}A}{{\rm d}\varrho}$. By a direct computation we get
$\frac{\partial\rho_{\mathrm{p}}^{\mathrm{st}}(\lambda,A)}{\partial A}=-A^{-1}\rho_{\mathrm{p}}^{\mathrm{st}}(\lambda)(1-\chi n^{\mathrm{st}}(\lambda)).$ (85)
@@
$\frac{\delta\mathcal{I}}{\delta\rho_{\mathrm{p}}}\Big|_{\rho_{\mathrm{p}}^{\mathrm{st}}}\cdot a=\frac{(\lambda-u)\rho_{\mathrm{p}}^{\mathrm{st}}(\lambda)(1-\chi n^{\mathrm{st}}(\lambda))}{\left(h_{0}|h_{0}\right)_{\mathrm{st}}}.$ (90)
@@
$\times\left[\delta\epsilon(\lambda)-\delta\epsilon(\lambda^{\prime})\right],$
where we used that for a stationary state equation (84) is satisfied. From the definition (82) it follows that $\delta\rho_{\mathrm{p}}=-\rho_{\mathrm{p}}(1-\chi n)\delta\epsilon$, which can be checked by a direct calculation. Introducing
$b(\lambda)=-\gamma\frac{\left(h_{0}|h_{0}\right)_{\mathrm{st}}\,a(\lambda)}{\rho_{\mathrm{p}}^{\mathrm{st}}(\lambda)(1-\chi n^{\mathrm{st}}(\lambda))}$ (92)
@@
Figure 4: Diffusion (panels (a) and (b)) and advective (panels (c),(d)) coefficients as a function of density $\varrho$ for the stochastic velocity swap models coupled to the classical bath. The bath is in the same state as in Fig. 5 and the collsion rate $\gamma=0.08$ is the same as well. Note that fermionic statstics make both diffusion and advective coefficients density-dependent. This has important consequences for the dynamics as visible in Fig. 5(d).
Figure 5: Molecular dynamics simulations for velocity swap models ($\gamma=0.08$, $\varrho^{(b)}=50.0$, $N=1600$) with initial states given by (104). We observe an excellent agreement between numerical simulations (histograms) and hydrodynamic equation (2) derived through ChE method (solid lines) for all cases. For symmetric states in bath system, cf. panels (a,b), both systems are described by diffusion equation (1). Upon introducing the asymmetry, the classical system is described by boosted diffusion equation, see panel (c). On the other hand, the fermionic model exhibits Burgers equation (2) dynamics, as visible in the dynamically generated asymmetry in the initially Gaussian shape, cf. panel (d). Results from microscopic simulations were averaged over 1000 realizations.
@@
For the classical particles $\chi=0$ and the relation between the bath state and a stationary state is especially simple and given by (20). The convection velocity function $u(\varrho)$ simplifies to
@@ VI Numerical simulations
$\rho_{\rm p}^{(b)}\propto e^{-(\lambda+\lambda_{B})^{2}/\sigma^{2}}+be^{-(\lambda-\lambda_{B})^{2}/\sigma^{2}}$ (104)
with $\sigma=6$, $\lambda_{B}=5.9$ and $b=1$ for symmetric states and $b=0.3$ for asymmetric states. The particles propagate freely and occasionally undergo two-body collisions with the bath. A collision amounts to velocity swap between system and a bath. When particles are classical and collisions happen at a constant rate the corresponding Master equation is precisely (38). Since the rate does not depend on the state of the system, such collisions can be efficiently implemented with standard Gillespie algorithm [82]. The more complicated case is the simulation of fermionic kinetic equation given by (39). There, we partition the system into fluid cells and propose an algorithm, which effectively incorporates the Pauli principle into the collisions. Algorithms for both cases are discussed in Appendix C.
The results are presented in Fig. 5. A very good agreement is found between molecular simulations and hydrodynamic equation (2). Depending on symmetry of the state (or lack thereof) and particle statistics, we find that numerical results excellently agree with diffusion equation (1) and with Burgers equation (2) for fermionic model with parity-asymmetric state.
Burgers equation in 1D is well-known for its breakdown of diffusive behavior in hydrodynamic two-point functions of the density field [83] close to homogenous equilibrium state. From this perspective, it is interesting to note that (finite) diffusion constant computed from kinetic theory is relevant for description of the numerics in our highly inhomogenous setting.
@@ VII Summary and discussion
In this work, we have considered a canonical problem of diffusion in which a minority species undergoes an irreversible motion due to coupling to the majority species, whose dynamics is not monitored and which, as is customary to assume, is in a spatially homogenous state. Due to the conservation of particle number of the minority species, its density obeys a continuity equation. In the standard situation of classical particles this continuity equation leads to a (possibly non-linear) diffusion equation. In this work we have identified the conditions under which the density follows instead the Burgers equation.
@@
The two sufficient conditions are: a) athermal and parity-asymmetric state of the majority species (bath) and b) quantum statistics of minority species (system).
@@
We have demonstrated that these two conditions are fulfilled for effective dynamics of nearly-integrable coupled gases. In such scenario, the integrability naturally makes athermal states possible. We have provided a kinetic explanation for appearance of such athermal states through limitations of two-body scattering in 1D. In our approach, we assume a large density imbalance between the two gases allowing for a separation of timescales between the slow dynamics in the dense tube (bath) and fast dynamics in the dilute tube (system). Starting from the GHD-Boltzmann description, we have derived an effective hydrodynamic equation for the density of the system particles (minority species). We have shown that the resulting equations become a standard diffusion equation when the bath is in a thermal state. Instead, for athermal state of a bath we obtain Burgers equation. Our method provides also expressions for transport coefficients: advective coefficient $u(\varrho)$ and diffusion constant $\mathcal{D}(\varrho)$.
@@
To unravel the essential ingredients leading to the Burgers equation we have considered also a simple stochastic model of two-particle scatterings between system and the bath. We observed an excellent agreement between microscopic numerical simulations and hydrodynamics derived from kinetic equation describing such random scattering events.
@@
Lastly, let us discuss the features of the hydrodynamic density-density correlation function, which is a natural object appearing in the literature on 1D Burgers equation [83]. This is because under the presence of ballistic mode with a nonlinear current, the correlation function is expected to exhibit a breakdown of diffusive behavior at the largest scales [24]. Close to the integrability, we generically expect large diffusion constants, which push away the anomalous broadening to very large space-time scales. In this sense, we expect a scenario similar to the one analyzed in the context of Navier-Stokes equations [60]. We also note that one has to be careful when looking at asymptotically large scales in models with (late-time) thermalization channels. There are three-particle (and higher) collision processes, which provide a mechanism for thermalization at late times, assuming that the density of the bath is finite. This leads ultimately to the breakdown of Burgers equation. We plan to investigate the interesting problem of two-point functions using the models studied here in the future.
@@ Acknowledgements.
Appendix A Grazing limit of the collision integral (19)
In this Appendix we formalize the grazing limit introduced in Sec. III. Our starting point is collision integral  (19). We introduce new variables under the integral
$\alpha=\lambda-\lambda^{\prime},\qquad\gamma=\frac{\lambda+\lambda^{\prime}}{2}$ (105)
and similarly we define $\alpha_{b}$ and $\gamma_{b}$. Note that the parameters $\alpha,\alpha_{b}$ are nothing else but changes of momenta of system (bath) particles that take part in a collision process $\lambda^{\prime}+\lambda_{b}^{\prime}\to\lambda+\lambda_{b}$. The collision rate becomes
$\mathcal{W}\left(\gamma+\frac{\alpha}{2},\gamma_{b}+\frac{\alpha_{b}}{2};\gamma-\frac{\alpha}{2},\gamma_{b}-\frac{\alpha_{b}}{2}\right)$ (106)
and the grazing limit is easiest to formulate when the expression above is viewed as a function of $\alpha$ and $\alpha_{b}$. We essentially assume that it is localized in the region where both $\alpha,\alpha_{b}$ are small as compared to typical range of variation of states $\rho_{\mathrm{p}}$ and $\rho_{\mathrm{p}}^{(b)}$. This allows for performing expansions in $\alpha,\alpha_{b}$ in other terms, that enter the integral. For convenience, we rewrite the collision integral into the form
$\mathcal{I}[\rho_{\mathrm{p}}](\mu)$ $=\int{\rm d}\lambda{\rm d}\lambda^{\prime}{\rm d}\lambda_{b}{\rm d}\lambda_{b}^{\prime}\delta(\mu-\lambda)\mathcal{W}(\lambda,\lambda_{b};\lambda^{\prime},\lambda_{b}^{\prime})\times$ (107)
$\delta(k)\delta(\omega)\rho_{\mathrm{p}}(\lambda^{\prime})\rho_{\mathrm{p}}^{(b)}(\lambda)\times$
$\left(\exp\left[\epsilon(\lambda)+\epsilon_{b}(\lambda_{b})-\epsilon(\lambda^{\prime})-\epsilon_{b}(\lambda_{b}^{\prime})\right]-1\right),$
For instance, the momentum conservation becomes
$k^{\prime}(\gamma)\alpha+k_{b}^{\prime}(\gamma_{b})\alpha_{b}=0$ (108)
and similarly for energy and expression with pseudoenergies. This yields
$\delta(k)\delta(\omega)=\frac{\delta\left(v(\gamma)-v_{b}(\gamma_{b})\right)\delta\left(\alpha_{b}+\alpha\frac{k^{\prime}(\gamma)}{k^{\prime}_{b}(\gamma_{b})}\right)}{|\alpha|k^{\prime}(\gamma)k_{b}^{\prime}(\gamma_{b})}.$ (109)
We also expand $\delta(\mu-\gamma-\frac{\alpha}{2})=\delta(\mu-\gamma)-\partial_{\mu}\delta(\mu-\gamma)\frac{\alpha}{2}+\ldots$. Moreover, we assume that the rate (106) is symmetric function of $\alpha$ and thus the first term in the expansion of Dirac delta vanishes. We define
$\tilde{\mathcal{W}}(\gamma,\gamma_{b})=\frac{1}{k^{\prime}(\gamma)k^{\prime}(\gamma_{b})}\int{\rm d}\alpha\frac{|\alpha|}{2}\times$ (110)
$\mathcal{W}\left(\gamma+\frac{\alpha}{2},\gamma_{b}-\alpha\frac{k^{\prime}(\gamma)}{k_{b}^{\prime}(\gamma_{b})};\gamma-\frac{\alpha}{2},\gamma_{b}+\alpha\frac{k^{\prime}(\gamma)}{k_{b}^{\prime}(\gamma_{b})}\right)$
and find that the collision integral indeed takes the form (21) upon truncating the Dirac delta expansion after the second term.
Appendix B Derivation of collision operator from GBBGKY hierarchy
The following section is based heavily on formalism and notation introduced in [48], especially on the part related to many-tube systems. Thus, for the purpose of clearer presentation in this section we change the notation slightly. The quasimomentum dependence will be denoted in lower index and we introduce rotation matrices $R=\mathbf{1}-\mathcal{T}n$. We also denote $\mathcal{A}_{\lambda_{1},\lambda_{2}}=R^{-t}_{\lambda_{1},\gamma}v_{\gamma}R^{t}_{\gamma,\lambda_{2}}$ and $\mathcal{C}_{\lambda_{1},\lambda_{2}}=R^{-t}_{\lambda_{1},\gamma}\rho_{\gamma}f_{\gamma}R^{-1}_{\gamma,\lambda_{2}}$. The upper indices $(i)$ denote tube index. To system tube we will refer to with index $(1)$, bath tube will be denoted with $(2)$. Einstein summation convention will be assumed in integration over repeated quasimomenta. We start with the following set of equations:
@@ Appendix B Derivation of collision operator from GBBGKY hierarchy
$g_{\lambda_{1},\lambda_{2}}=(R^{(1)})^{-t}_{\lambda_{1},\gamma_{1}}(R^{(2)})^{-t}_{\lambda_{2},\gamma_{2}}\bar{g}_{\gamma_{1},\gamma_{2}},$ (116)
@@
$\mathcal{G}_{\lambda_{1},\lambda_{2}}(z,z^{\prime};t,t^{\prime})=\frac{1}{2\pi}\int{\rm d}k\,e^{ik(z+z^{\prime}-v_{\lambda_{1},\lambda_{2}}t)}\theta_{H}(t-t^{\prime})$ (118)
@@
$\bar{g}_{\lambda_{1},\lambda_{2}}(z,t)=\int{\rm d}z^{\prime}{\rm d}t^{\prime}\mathcal{G}_{\lambda_{1},\lambda_{2}}(z,z^{\prime};t,t^{\prime})\bar{\mathcal{S}}_{\lambda_{1},\lambda_{2}}(z^{\prime},t^{\prime}).$ (119)
@@
In the next step we plug this expression into (113), and perform the standard approximations used in such derivations [75], related to separations of scales. In practice replace the time argument of $\mathcal{S}_{\eta_{1},\eta_{2}}[\rho](z^{\prime},t-\tau)$ with $t$ and extend the integration range over $\tau$ to infinity. We also insert momentum representations of interaction potential $V(z)=\int{\rm d}ke^{ikz}\tilde{V}(k)$ and use the property that $\int{\rm d}\lambda R^{-t}_{\lambda,\gamma}=2\pi\rho^{\rm t}_{\gamma}$. After all these steps we find the collision integral
$\partial_{t}\rho_{\lambda_{1}}^{(1)}=\mathcal{I}(\lambda_{1}),\qquad\mathcal{I}(\lambda_{1})=\partial_{\lambda_{1}}\left[(R^{(1)})^{-t}_{\lambda_{1},\eta}\mathcal{I}_{0}(\eta)\right],$ (121)
@@
$\mathcal{I}_{0}(\lambda_{1})=(2\pi)^{2}\int{\rm d}kk^{2}\tilde{V}^{2}(k)\int_{0}^{\infty}{\rm d}\tau\int{\rm d}\lambda_{2}e^{-ikv_{\lambda_{1}\lambda_{2}}\tau}\rho^{\rm t,(2)}_{\lambda_{2}}(R^{(1)})_{\lambda_{1},\eta_{1}}^{t}(R^{(2)})_{\lambda_{2},\eta_{2}}^{t}\int{\rm d}\eta_{3}\left(\partial_{\eta_{1}}\rho_{\eta_{1}}^{(1)}\mathcal{C}^{(2)}_{\eta_{2}\eta_{3}}-\partial_{\eta_{2}}\rho_{\eta_{2}}^{(2)}\mathcal{C}^{(1)}_{\eta_{1}\eta_{3}}\right)$. (122)
Where by $\rho^{\rm t,(2)}_{\lambda_{2}}$ we denoted total density of states in tube 2. We proceed by using [75] $\int_{0}^{\infty}{\rm d}\tau e^{i\tau x}=\pi\delta(x)+i\mathcal{P}(\frac{1}{x})$, where the second term is the principal value, which vanishes as the integrand is a symmetric function in $k$. In what follows we also use that [48] $\int{\rm d}\lambda_{2}\,\mathcal{C}_{\lambda_{1},\lambda_{2}}=2\pi R_{\lambda_{1},\gamma}^{-t}\rho_{\gamma}f_{\gamma}\rho^{\rm t}_{\gamma}$ and $R^{t}_{\lambda,\gamma}\partial_{\gamma}\rho_{\gamma}=-\rho_{\lambda}f_{\lambda}\partial_{\lambda}\epsilon_{\lambda}$. We get
$\mathcal{I}_{0}(\lambda_{1})=-2\pi^{2}\int{\rm d}k|k|\tilde{V}^{2}(k)\int{\rm d}\lambda_{2}\delta(v_{\lambda_{1},\lambda_{2}})\left(k_{2}^{\prime}(\lambda_{2})\right)^{2}\rho_{\lambda_{1}}^{(1)}f_{\lambda_{1}}^{(1)}\rho_{\lambda_{2}}^{(2)}f_{\lambda_{2}}^{(2)}\left(\epsilon^{\prime}_{\lambda_{1}}-\frac{k^{\prime}_{1}(\lambda_{1})}{k^{\prime}_{2}(\lambda_{2})}\epsilon^{\prime}_{\lambda_{2}}\right).$ (123)
@@
Appendix C Molecular dynamics simulation algorithm
The classical system with collision integral given by (38) can be simulated fairly easy with a standard implementation of Gillespie algorithm [82]. The fermionic case (39) is more complicated due to Fermi exclusion principle, which we want to somehow model using dynamics of classical particles. We thus wish to construct a classical model for which the dynamics in phase-space would be described by kinetic equation (39). Our approach will turn out to be more general and cover both classical and fermionic case.